# On the Robustness of Kernel Goodness-of-Fit Tests

### v4

@@ 2.3 Robustness for Goodness-of-Fit Testing
Robustness of GOF tests refers to the lack of sensitivity of the test outcome to small model deviations (Rieder,, 1982; Lambert,, 1982). Model deviations can be formalized as a neighborhood $\mathcal{P}_{0}$ around a nominal model $P$. The neighborhood $\mathcal{P}_{0}$ is often called an uncertainty set, and encodes the practitioner’s uncertainty on the model $P$. One popular example for $\mathcal{P}_{0}$ is Huber’s contamination model (Huber,, 1964, 1965)
@@
where $\epsilon\in[0,1]$ is the maximal contamination ratio, and the probability measure $R$ acts as arbitrary contamination. This model is appropriate when practitioners believe that $P$ accurately describes all but a small proportion of the data. Huber’s models have been studied in the context of robust GOF testing including Huber, (1965); Qin and Priebe, (2017), as well as in robust estimation (Hampel,, 1974; Huber and Ronchetti,, 2011). Special interests lie in the case when $R$ is restricted to point masses, i.e., $R=\delta_{\mathbf{z}}$ for some $\mathbf{z}\in\mathbb{R}^{d}$, where robustness is often called bias-robustness (Huber and Ronchetti,, 2011). Beyond Huber’s models, $\mathcal{P}_{0}$ can also be chosen as density-band models (Kassam,, 1981), which assume the density of the data-generating distribution is close to the model density up to a small error (see Section˜4.2). Moreover, $\mathcal{P}_{0}$ can be set to a ball defined via a statistical divergence or metric, such as Hellinger distance (LeCam,, 1973), Wasserstein metric (Gao et al.,, 2018), and Maximum Mean Discrepancy (Sun and Zou,, 2023).
Uncertainty sets $\mathcal{P}_{0}$ provide a framework for assessing the robustness of a GOF test targeting the classic null $H_{0}:Q=P$. By studying the probability of rejection when the data-generating distribution $Q$ deviates from the model $P$ but remains within $\mathcal{P}_{0}$, we can gauge the test’s robustness. In the robust testing literature, a common approach is to consider a sequence $\{\mathcal{P}_{0}^{n}\}_{n=1}^{\infty}$ of uncertainty sets with shrinking size as the sample size $n$ increases. This is formalized in the following notion of qualitative robustness, inspired by Rieder, (1982).
@@
Let $\mathcal{P}_{0}^{1}\supseteq\mathcal{P}_{0}^{2}\supseteq\ldots$ be a sequence of subsets of $\mathcal{P}(\mathbb{R}^{d})$ that contain $P$ and such that $\cap_{n=1}^{\infty}\mathcal{P}_{0}^{n}=\{P\}$. For any positive integer $n$, let $T_{n}:(\mathbb{R}^{d})^{n}\to\mathbb{R}$ be a test statistic where a large value of $T_{n}$ suggests deviation from the null $H_{0}:Q=P$, and let $\gamma_{n}:(\mathbb{R}^{d})^{n}\to\mathbb{R}$ be a function that computes the decision threshold. A sequence of hypothesis tests (indexed by $n$) that reject $H_{0}$ when $T_{n}>\gamma_{n}$ is qualitatively robust to $\{\mathcal{P}_{0}^{n}\}_{n=1}^{\infty}$ if, as $n\to\infty$,
@@
Intuitively, this notion of robustness asks how sensitive the test outcome is under infinitesimally small model deviation. The shrinking-size condition ensures that the probability of rejection under any $Q\neq P$ does not trivially approach one due to the consistency of the test. By choosing the $\mathcal{P}_{0}^{n}$ in Definition˜1 to be Prokhorov balls (Prokhorov,, 1956), we would recover the qualitative robustness introduced in Rieder, (1982, Definition 2.1), which parallels the conventional notion of qualitative robustness for estimators as argued in Rieder, (1982, Remark 2). However, distributions in Prokhorov balls do not have a simple form, posing challenges to the analysis. Our definition extends it to a general sequence of neighborhood. In Section˜3, we will choose $\mathcal{P}_{0}^{n}$ to be Huber’s contamination models, which both significantly simplifies the analysis and encompasses a wide range of relevant scenarios. We will show that, within Huber’s neighborhood, the standard KSD test is not qualitatively robust with stationary kernels, but is qualitatively robust with appropriately tilted kernels.
Moreover, our definition of qualitative robustness is tied to a sequence of shrinking neighborhood $\{P_{0}^{n}\}_{n=1}^{\infty}$. Clearly, if a test is qualitatively robust to a sequence $\{P_{0}^{n}\}_{n=1}^{\infty}$, then it is also qualitatively robust to any sequences that decay faster. Therefore, loosely speaking, the rate of decay of $\{P_{0}^{n}\}_{n=1}^{\infty}$ characterizes the degree of qualitative robustness. In Section˜3, we will explicitly derive the rate required for the standard KSD test to maintain qualitative robustness.
However, qualitative robustness has its own limitations. It only concerns the insensitivity of a test to sufficiently small model deviations, but offers no guarantees for deviations of a fixed size. The latter scenario is more practically pertinent, as practitioners often need to account for a specific form of model misspecification and require the test to remain well-calibrated under that level of uncertainty. This can be formalized by relaxing the point null hypothesis $H_{0}$ to a composite hypothesis $H^{\mathrm{C}}_{0}:Q\in\mathcal{P}_{0}$ for some uncertainty set $\mathcal{P}_{0}$, and requiring calibration under $H^{\mathrm{C}}_{0}$. This motivates the notion of quantitative robustness.
Definition 2 (Quantitative robustness to a neighborhood).
@@
Quantitatively robust GOF tests have been developed for various types of uncertainty sets $\mathcal{P}_{0}$, including Huber’s contamination model (Huber,, 1965) and neighborhoods defined by Kullback-Leibler divergence (Levy,, 2008; Yang and Chen,, 2018) or $\alpha$-divergence (Gül and Zoubir,, 2016). These tests enjoy minimax optimality but require the normalizing constant of $P$ to be known, thus not applicable to unnormalized models. Two-sample tests that are quantitatively robust to Hellinger distance (LeCam,, 1973), Wasserstein distance (Gao et al.,, 2018), and Maximum Mean Discrepancy (Sun and Zou,, 2023; Gao et al.,, 2021) have also been proposed. However, to be used for GOF testing, they require approximation of $P$ by finite samples, which incurs extra approximation error and can be computationally demanding due to the non-trivial task of sampling from $P$.
In Section˜4, we will choose $\mathcal{P}_{0}$ to be a KSD ball centered at $P$. We choose KSD balls over neighborhood defined by other statistical divergences because (i) this choice naturally lends itself to a GOF test that is both easy to implement and guarantees robustness at little extra computational cost, and (ii) it is easy to select the radius of such balls so that our proposed test is quantitatively robust to various types of contamination of interest, such as Huber’s contaminations or density-band contaminations.
Qualitative versus quantitative robustness
Qualitative robustness concerns the limiting behavior of a test against some specific class of local alternatives. Thus, it is closely tied to the minimax separation boundary of a test (Ingster,, 1987, 1993). Moreover, qualitative robustness offers a notion of “degree of robustness” of a given test via the decay rate of the uncertainty neighbouhood sequence $\{\mathcal{P}_{0}^{n}\}_{n=1}^{\infty}$. In contrast, quantitative robustness cannot be used for this purpose, because any consistent test is not quantitatively robust by definition. However, quantitative robustness has more practical relevance, as it concerns a fixed composite null set, which can be explicitly constructed to enforce robustness. This is why we have introduced two different notions of robustness.
@@ 3 The (Lack of) Robustness of Existing Kernel Goodness-of-fit Tests
We will now study the robustness of standard KSD GOF tests using stationary kernels in Section˜3.1, then study their tilted counterparts which are popular in the parameter estimation literature in Section˜3.2.
@@ 3.1 Existing KSD Tests with Stationary Kernels are not Qualitatively Robust
$\sup_{Q\in\mathcal{P}(P;\epsilon_{n})}\big{|}\Pr\nolimits_{\mathbb{X}_{n}\sim Q,\mathbf{W}}\big{(}D^{2}(\mathbb{X}_{n})>q^{2}_{\infty,1-\alpha}(\mathbb{X}_{n})\big{)}-\Pr\nolimits_{\mathbb{X}_{n}^{\ast}\sim P,\mathbf{W}}\big{(}D^{2}(\mathbb{X}_{n}^{\ast})>q^{2}_{\infty,1-\alpha}(\mathbb{X}_{n}^{\ast})\big{)}\big{|}\;$  $\to\;1-\alpha\;,$
@@
Remark 1.1 (Unbounded Stein kernel).
Theorem˜1 is a consequence of the fact that the Stein kernel evaluated at $\mathbf{x}$, i.e., $\mathbf{x}\mapsto u_{p}(\mathbf{x},\mathbf{x})$, is unbounded when $P$ has an exploding score function. Exploding score functions are often associated with light tails. Examples of such distributions include those with a density of the form $p(\mathbf{x})\propto\exp(-\|\mathbf{x}\|_{2}^{r})$ with $r\geq 2$. These distributions have sub-Gaussian tails, and their score function has the form $\mathbf{s}_{p}(\mathbf{x})=-r\mathbf{x}\|\mathbf{x}\|_{2}^{r-2}$, which is unbounded for $r\geq 2$. On the other hand, this exploding-score condition excludes heavy-tailed models with bounded score functions, such as super-Laplacian or t-distributions (Gorham and Mackey,, 2017; Barp et al.,, 2024).
Remark 1.2 (Moment conditions).
The moment conditions in Theorem˜1 are mild. For the example from Remark˜1.1, namely $p(\mathbf{x})\propto\exp(-\|\mathbf{x}\|_{2}^{r})$ with $r\geq 2$, direct computation shows that these conditions hold if $h(\mathbf{x})$ and $\nabla h(\mathbf{x})$ decay at least as fast as $\||\mathbf{x}\|_{2}^{-(r-1)}$ so as to cancel the growth of the score $\|\mathbf{s}_{p}(\mathbf{x})\|_{2}=r\|\mathbf{x}\|_{2}^{r-1}$. Such kernels include squared-exponential kernels, IMQ kernels $h(u)=(1+\|u\|_{2}^{2})^{-s/2}$ with $s\geq r-1$, and Matérn kernels with sufficient smoothness. Section˜5.1 will provide numerical evidence using a Gaussian model.
Remark 1.3 (Connection to separation boundaries).
Let $\mathcal{Q}_{\delta}\coloneqq\{Q\in\mathcal{P}(\mathbb{R}^{d}):S(Q,P)\geq\delta\}$, where $S$ is some statistical divergence or metric. The separation boundary (Ingster,, 1987, 1993) of a test is the fastest decaying sequence $\{\delta_{n}\}_{n}$ such that the test power under any $Q\in\mathcal{Q}_{\delta_{n}}$ still converges to 1 as $n\to\infty$. From this perspective, Theorem˜1 implies that the separation boundary of the standard KSD test must decay at least with rate $\epsilon_{n}=o(n)^{-1}$, whenever $Q_{\delta_{n}}$ contains Huber-contamination models. This complements existing results on the separation boundary of KSD tests. The most relevant works are Schrab et al., (2022); Hagrass et al., (2025), where they consider alternative distributions $Q$ that have a density with respect to either the Lebesgue measure or the target model $P$. In particular, their results cannot be used to derive Theorem˜1, since their density assumption on $Q$ is violated in our case. This is because we consider alternatives of the form $Q=(1-\epsilon_{n})P+\epsilon_{n}\delta_{\mathbf{z}}$, which involve a Dirac delta measure and thus has no density with respect to the Lebesgue measure nor any continuous measures, including $P$.
@@ 3.2 Existing KSD Tests with Tilted Kernels are Qualitatively Robust
Since the lack of robustness of existing KSD tests with stationary kernels arises because the Stein kernel is unbounded, a natural approach is to choose another kernel $k$ which prevents this from happening. We now show that this can be achieved using tilted kernels, a common approach to induce robustness for parameter estimation (Barp et al.,, 2019; Matsubara et al.,, 2022). Specifically, we will show in Theorem˜3 that, with a tilted kernel, a small proportion of contamination in the data has negligible impact on the outcome of the standard KSD test.
We first show that tilted kernels give rise to bounded Stein kernels. The proof is deferred to Section˜A.2.
@@
Notably, the function $\mathbf{x}\mapsto u_{p}(\mathbf{x},\mathbf{x})$ is non-negative since $u_{p}$ is a reproducing kernel. Moreover, Barp et al., (2024, Theorem 9) shows that KSD with such tilted kernels still satisfies $P$-separation, namely $D(Q,P)=0\iff Q=P$, provided that $\|\mathbf{s}_{p}\|_{2}$ grows at most root-exponentially and the spectral density of the translation-invariant kernel kk—which exists by Bochner’s Theorem (Berlinet and Thomas-Agnan,, 2004, Theorem 20)—is bounded away from zero on compact sets.
Most stationary kernels used in KSD tests (such as the IMQ or squared-exponential kernels) satisfy the conditions in Lemma˜2. In this construction, the weighting function $w$ is used to counteract the growth of the score function $\mathbf{s}_{p}$. When the score grows polynomially like $\|\mathbf{s}_{p}(\mathbf{x})\|_{2}=\mathcal{O}(\|\mathbf{x}\|_{2}^{r})$ for some $r>0$, it suffices to choose $w(\mathbf{x})=(1+a^{2}\|\mathbf{x}\|_{2}^{2})^{-b}$ for any $a>0$ and $b\geq 1/(2r)$. This weight function is common in the frequentist parameter estimation literature (Barp et al.,, 2019) and the generalized Bayesian inference literature (Matsubara et al.,, 2022; Altamirano et al.,, 2023, 2024; Duran-Martin et al.,, 2024).
@@
Remark 3.2.
The boundedness condition on $\|w(\mathbf{x})\mathbf{s}_{p}(\mathbf{x})\|_{2}$ required of Lemma˜2 and Theorem˜3 bears similarity with the ones imposed for KSD-based robust estimation methods (Barp et al.,, 2024; Matsubara et al.,, 2022). For example, Barp et al., (2019, Proposition 7) showed that their KSD-based estimator is globally bias-robust assuming $\mathbf{x}\mapsto\|\mathbf{s}_{p}(\mathbf{x})\|_{2}\int_{\mathbb{R}^{d}}\|k(\mathbf{x},\mathbf{x}^{\prime})\mathbf{s}_{p}(\mathbf{x}^{\prime})\|_{2}Q(\mathop{}\!\mathrm{d}\mathbf{x}^{\prime})$ is bounded. This is slightly weaker than those we assume in Lemma˜2 and Theorem˜3, but also harder to verify in practice.
Remark 3.3 (Connection to separation boundaries, continued).
Theorem˜3 implies that the separation boundary of the KSD tests cannot decay faster than $n^{-1/2}$ when considering alternatives that include Huber-contamination models (see also Remark˜1.3). This can be compared to existing results on the separation boundaries of tests based on Maximum Mean Discrepancy (MMD, Müller,, 1997; Gretton et al.,, 2012). MMD is a broader family of discrepancies that recovers KSD when the kernel is chosen to be the Stein kernel $u_{p}$ (Barp et al.,, 2024, Theorem 1). The separation rates of MMD tests have been studied extensively (Balasubramanian et al.,, 2021; Hagrass et al., 2024b, ), but these works assume the alternative $Q$ has a density with respect to either the Lebesgue measure or $P$, which does not hold in our contamination setting. An exception is Hagrass et al., 2024a , which does not make this assumption. However, their results are limited to the two-sample setting, where the target distribution $P$ is unknown and approximated by finite samples, again differing from our setup.
Although tilted kernels enforce qualitative robustness, they are not enough to guarantee quantitative robustness. Indeed, so long as $k$ is characteristic, the consistency of the standard KSD test implies that $\Pr_{\mathbb{X}_{n}\sim Q,\mathbf{W}}(D^{2}(\mathbb{X}_{n})>q^{2}_{\infty,1-\alpha}(\mathbb{X}_{n}))\to 1$ whenever $Q\neq P$ Liu et al., 2016, Proposition 4.2; Schrab et al., 2022, Theorem 3.3, so KSD tests with either stationary or tilted kernels are not quantitatively robust to any neighborhoods strictly larger than the singleton set $\{P\}$. We provide numerical evidence in Section˜5.1.

### v5

@@ 2.3 Robustness for Goodness-of-Fit Testing
Robustness of GOF tests refers to the lack of sensitivity of the test outcome to small model deviations (Rieder, 1982; Lambert, 1982). Model deviations can be formalized as a neighborhood $\mathcal{P}_{0}$ around a nominal distribution $P$. The neighborhood $\mathcal{P}_{0}$ is often called an uncertainty set, and encodes the practitioner’s uncertainty on $P$. One popular example for $\mathcal{P}_{0}$ is Huber’s contamination model (Huber, 1964, 1965)
@@
where $\epsilon\in[0,1]$ is the maximal contamination ratio, and the probability measure $R$ acts as arbitrary contamination. This model is appropriate when practitioners believe that $P$ accurately describes all but a small proportion of the data.
Huber’s models have been studied in the context of robust GOF testing including Huber (1965); Qin and Priebe (2017), as well as in robust estimation (Hampel, 1974; Huber and Ronchetti, 2011). Special interests lie in the case when $R$ is restricted to point masses, i.e., $R=\delta_{\mathbf{z}}$ for some $\mathbf{z}\in\mathbb{R}^{d}$, where robustness is often called bias-robustness (Huber and Ronchetti, 2011). Beyond Huber’s models, $\mathcal{P}_{0}$ can also be chosen as density-band models (Kassam, 1981), which assume the density of the data-generating distribution is close to the model density up to a small error (see Section˜4.2). Moreover, $\mathcal{P}_{0}$ can be set to a ball defined via a statistical divergence or metric, such as Hellinger distance (LeCam, 1973), Wasserstein metric (Gao et al., 2018), and Maximum Mean Discrepancy (Sun and Zou, 2023).
Uncertainty sets $\mathcal{P}_{0}$ provide a framework for assessing the robustness of a GOF test, by studying the rejection probability when $Q$ deviates from the nominal distribution $P$ but remains within $\mathcal{P}_{0}$. In the robust testing literature, a common approach is to consider a sequence $\{\mathcal{P}_{0}^{n}\}_{n=1}^{\infty}$ of uncertainty sets with shrinking size as the sample size $n$ increases. This is formalized in the following notion of qualitative robustness, inspired by Rieder (1982).
@@
Let $\mathcal{P}_{0}^{1}\supseteq\mathcal{P}_{0}^{2}\supseteq\ldots$ be a sequence of subsets of $\mathcal{P}(\mathbb{R}^{d})$ that contain $P$ and such that $\cap_{n=1}^{\infty}\mathcal{P}_{0}^{n}=\{P\}$. For any positive integer $n$, let $T_{n}:\big(\mathbb{R}^{d}\big)^{n}\to\mathbb{R}$ be a test statistic where a large value of $T_{n}$ suggests deviation from the null $H_{0}:Q=P$, and let $\gamma_{n}:\big(\mathbb{R}^{d}\big)^{n}\to\mathbb{R}$ be a function that computes the decision threshold. A sequence of hypothesis tests (indexed by $n$) that reject $H_{0}$ when $T_{n}>\gamma_{n}$ is qualitatively robust to $\{\mathcal{P}_{0}^{n}\}_{n=1}^{\infty}$ if, as $n\to\infty$,
@@
Intuitively, this notion of robustness asks how sensitive the test outcome is under infinitesimally small model deviation. The shrinking-size condition ensures that the rejection probability under any $Q\neq P$ does not trivially approach one due to the consistency of the test. By choosing the $\mathcal{P}_{0}^{n}$ in Definition˜1 to be Prokhorov balls (Prokhorov, 1956), we would recover the qualitative robustness introduced in Rieder (1982, Definition 2.1), which parallels the conventional notion of qualitative robustness for estimators (Rieder, 1982, Remark 2). However, distributions in Prokhorov balls do not have a simple form, posing challenges to the analysis. Our definition extends it to a general sequence of neighborhood. In Section˜3, we will choose $\mathcal{P}_{0}^{n}$ to be Huber’s contamination models (6), which both significantly simplifies the analysis and encompasses a wide range of relevant scenarios. We will show that, within Huber’s neighborhood, the standard KSD test is not necessarily qualitatively robust with stationary kernels, but is qualitatively robust with appropriately tilted kernels.
Moreover, our definition of qualitative robustness is tied to a sequence of shrinking neighborhood $\{P_{0}^{n}\}_{n=1}^{\infty}$. Clearly, if a test is qualitatively robust to a sequence $\{P_{0}^{n}\}_{n=1}^{\infty}$, then it is also qualitatively robust to any sequences that decay faster. Therefore, loosely speaking, the rate of decay of $\{P_{0}^{n}\}_{n=1}^{\infty}$ characterizes the degree of qualitative robustness. In Section˜3, we will explicitly derive the rate required for the standard KSD test to retain qualitative robustness.
However, qualitative robustness has its own limitations. It only concerns the insensitivity of a test to sufficiently small model deviations, but offers no guarantees for deviations of a fixed size. The latter scenario is more practically pertinent, as practitioners often need to account for a specific form of model misspecification and require the test to remain well-calibrated under that level of uncertainty. This can be formalized by relaxing the point null hypothesis $H_{0}$ to a composite hypothesis $H^{\mathrm{C}}_{0}:Q\in\mathcal{P}_{0}$, and requiring calibration under $H^{\mathrm{C}}_{0}$. This motivates the following notion of quantitative robustness.
Definition 2 (Quantitative robustness to a single neighborhood).
@@
Quantitatively robust GOF tests have been developed for various types of uncertainty sets $\mathcal{P}_{0}$, including Huber’s contamination model (Huber, 1965) and neighborhoods defined by Kullback-Leibler divergence (Levy, 2008; Yang and Chen, 2018) or $\alpha$-divergence (Gül and Zoubir, 2016). These tests enjoy minimax optimality but require the normalizing constant of $P$ to be known, thus not applicable to unnormalized models. Two-sample tests that are quantitatively robust to Hellinger distance (LeCam, 1973), Wasserstein distance (Gao et al., 2018), and Maximum Mean Discrepancy (Sun and Zou, 2023; Gao et al., 2021) have also been proposed. However, to be used for GOF testing, they require approximation of $P$ by finite samples, which incurs extra approximation error and can be computationally demanding due to the non-trivial task of sampling from $P$.
In Section˜4, we will choose $\mathcal{P}_{0}$ to be a KSD ball centered at $P$. We choose KSD balls over other types of neighborhood because (i) this choice naturally lends itself to a GOF test that is both easy to implement and guarantees robustness at little extra computational cost, and (ii) it is easy to select the radius of such balls so that our proposed test is quantitatively robust to various types of contamination of interest, such as Huber’s contaminations or density-band contaminations.
We conclude this section with a brief comparison between qualitative and quantitative robustness. Qualitative robustness concerns the limiting behavior of a test against some specific class of local alternatives. Thus, it is closely tied to the minimax separation boundary of a test (Ingster, 1987, 1993). Moreover, it offers a notion of “degree of robustness” of a given test via the decay rate of the uncertainty neighbouhood sequence $\{\mathcal{P}_{0}^{n}\}_{n=1}^{\infty}$. In contrast, quantitative robustness cannot be used for this purpose, because any consistent test is not quantitatively robust by definition. However, quantitative robustness has more practical relevance, as it concerns a fixed composite null set, which can be explicitly constructed to enforce robustness. This is why we have introduced two different notions of robustness.
@@ 3 The (Lack of) Robustness of Existing Kernel Goodness-of-fit Tests
We will now study the robustness of standard KSD GOF tests using stationary kernels in Section˜3.1, and then their tilted counterparts which are popular in the parameter estimation literature in Section˜3.2.
@@ 3.1 Existing KSD Tests with Stationary Kernels are not Qualitatively Robust
$\sup_{Q\in\mathcal{P}(P;\epsilon_{n})}\big|\Pr\nolimits_{\mathbb{X}_{n}\sim Q,\mathbf{W}}\big(D^{2}(\mathbb{X}_{n})>q^{2}_{\infty,1-\alpha}(\mathbb{X}_{n})\big)-\Pr\nolimits_{\mathbb{X}_{n}^{\ast}\sim P,\mathbf{W}}\big(D^{2}(\mathbb{X}_{n}^{\ast})>q^{2}_{\infty,1-\alpha}(\mathbb{X}_{n}^{\ast})\big)\big|$
 $\to\;1-\alpha\;,$
@@
Remark 1 (Unbounded Stein kernel).
Theorem˜1 is a consequence of the fact that the Stein kernel evaluated at $\mathbf{x}$, i.e., $\mathbf{x}\mapsto u_{p}(\mathbf{x},\mathbf{x})$, is unbounded when $P$ has an exploding score function. Exploding score functions are often associated with light tails. Examples of such distributions include those with a density of the form $p(\mathbf{x})\propto\exp(-\|\mathbf{x}\|_{2}^{r})$ with $r\geq 2$. These distributions have sub-Gaussian tails, and their score function has the form $\mathbf{s}_{p}(\mathbf{x})=-r\mathbf{x}\|\mathbf{x}\|_{2}^{r-2}$, which is unbounded for $r\geq 2$. On the other hand, this exploding-score condition excludes heavy-tailed models with bounded score functions, such as super-Laplacian or t-distributions (Gorham and Mackey, 2017; Barp et al., 2024).
Remark 2 (Moment conditions).
The moment conditions in Theorem˜1 are mild. For the example from Remark˜1, namely $p(\mathbf{x})\propto\exp(-\|\mathbf{x}\|_{2}^{r})$ with $r\geq 2$, direct computation shows that these conditions hold if $h(\mathbf{x})$ and $\nabla h(\mathbf{x})$ decay at least as fast as $\||\mathbf{x}\|_{2}^{-(r-1)}$ so as to cancel the growth of the score $\|\mathbf{s}_{p}(\mathbf{x})\|_{2}=r\|\mathbf{x}\|_{2}^{r-1}$. Such kernels include squared-exponential kernels, IMQ kernels $h(u)=(1+\|u\|_{2}^{2})^{-s/2}$ with $s\geq r-1$, and Matérn kernels with sufficient smoothness. Section˜5.1 will provide numerical evidence using a Gaussian model.
Remark 3 (Connection to separation boundaries).
Let $\mathcal{Q}_{\delta}\coloneqq\{Q\in\mathcal{P}(\mathbb{R}^{d}):S(Q,P)\geq\delta\}$, where $S$ is some statistical divergence or metric. The separation boundary (Ingster, 1987, 1993) of a test is the fastest decaying sequence $\{\delta_{n}\}_{n}$ such that the test power under any $Q\in\mathcal{Q}_{\delta_{n}}$ still converges to 1 as $n\to\infty$. From this perspective, Theorem˜1 implies that the separation boundary of the standard KSD test must decay at least with rate $\epsilon_{n}=o(n)^{-1}$, whenever $Q_{\delta_{n}}$ contains Huber-contamination models. This complements existing results on the separation boundary of KSD tests. The most relevant works are Schrab et al. (2022); Hagrass et al. (2025), where they consider alternative distributions $Q$ that have a density with respect to either the Lebesgue measure or the target model $P$. In particular, their results cannot be used to derive Theorem˜1, since their density assumption on $Q$ is violated in our case. Indeed, we consider alternatives of the form $Q=(1-\epsilon_{n})P+\epsilon_{n}\delta_{\mathbf{z}}$, which involve a Dirac delta measure and thus has no density with respect to the Lebesgue measure nor any continuous measures, including $P$.
@@ 3.2 Existing KSD Tests with Tilted Kernels are Qualitatively Robust
Since unbounded Stein kernels are the main cause of the lack of robustness of existing KSD tests, a natural approach to enforce robustness is to choose a suitable $k$ so that the Stein kernel $u_{p}$ becomes bounded. We now show that this can be achieved using tilted kernels (Barp et al., 2019; Matsubara et al., 2022). Specifically, we will show in Theorem˜3 that, with a tilted kernel, a small proportion of contamination in the data has negligible impact on the outcome of the standard KSD test.
We first prove that tilted kernels give rise to bounded Stein kernels. The proof is deferred to Section˜A.2.
@@
Notably, the function $\mathbf{x}\mapsto u_{p}(\mathbf{x},\mathbf{x})$ is non-negative since $u_{p}$ is a reproducing kernel. Barp et al. (2024, Theorem 9) showed that KSD with such tilted kernels still satisfies $P$-separation, namely $D(Q,P)=0\iff Q=P$, provided that $\|\mathbf{s}_{p}\|_{2}$ grows at most root-exponentially and the spectral density of the translation-invariant kernel kk—which exists by Bochner’s Theorem (Berlinet and Thomas-Agnan, 2004, Theorem 20)—is bounded away from zero on compact sets.
Most stationary kernels used in KSD tests (such as the IMQ or squared-exponential kernels) satisfy the conditions in Lemma˜2. Intuitively, the weighting function $w$ is used to counteract the growth of the score function $\mathbf{s}_{p}$. When the score grows polynomially like $\|\mathbf{s}_{p}(\mathbf{x})\|_{2}=\mathcal{O}(\|\mathbf{x}\|_{2}^{r})$ for some $r>0$, it suffices to choose $w(\mathbf{x})=(1+a^{2}\|\mathbf{x}\|_{2}^{2})^{-b}$ for any $a>0$ and $b\geq 1/(2r)$. This weight function is common in the frequentist parameter estimation literature (Barp et al., 2019) and the generalized Bayesian inference literature (Matsubara et al., 2022; Altamirano et al., 2023, 2024; Duran-Martin et al., 2024).
@@
Remark 5.
The boundedness condition on $\|w(\mathbf{x})\mathbf{s}_{p}(\mathbf{x})\|_{2}$ required of Lemma˜2 and Theorem˜3 bears similarity with the ones imposed for KSD-based robust estimation methods (Barp et al., 2024; Matsubara et al., 2022). For example, Barp et al. (2019, Proposition 7) showed that their KSD-based estimator is globally bias-robust assuming $\mathbf{x}\mapsto\|\mathbf{s}_{p}(\mathbf{x})\|_{2}\int_{\mathbb{R}^{d}}\|k(\mathbf{x},\mathbf{x}^{\prime})\mathbf{s}_{p}(\mathbf{x}^{\prime})\|_{2}Q(\mathop{}\!\mathrm{d}\mathbf{x}^{\prime})$ is bounded. This is slightly weaker than those we assume in Lemma˜2 and Theorem˜3, but also harder to verify in practice.
Remark 6 (Connection to separation boundaries, continued).
Theorem˜3 implies that the separation boundary of the KSD tests cannot decay faster than $n^{-1/2}$ when considering alternatives that include Huber-contamination models (see also Remark˜3). This can be compared to existing results on the separation boundaries of tests based on Maximum Mean Discrepancy (MMD, Müller, 1997; Gretton et al., 2012). MMD is a broader family of discrepancies that recovers KSD when the kernel is chosen to be the Stein kernel $u_{p}$ (Barp et al., 2024, Theorem 1). The separation rates of MMD tests have been studied extensively (Balasubramanian et al., 2021; Hagrass et al., 2024b), but these works assume the alternative $Q$ has a density with respect to either the Lebesgue measure or $P$, which does not hold in our contamination setting. An exception is Hagrass et al. (2024a), which does not make this assumption. However, their results are limited to the two-sample setting, where the target distribution $P$ is unknown and approximated by finite samples, again differing from our setup.
Although tilted kernels enforce qualitative robustness, they are not enough to guarantee quantitative robustness. Indeed, so long as $k$ is characteristic, the consistency of the standard KSD test implies that $\Pr_{\mathbb{X}_{n}\sim Q,\mathbf{W}}(D^{2}(\mathbb{X}_{n})>q^{2}_{\infty,1-\alpha}(\mathbb{X}_{n}))\to 1$ whenever $Q\neq P$ Liu et al. 2016, Proposition 4.2; Schrab et al. 2022, Theorem 3.3, so KSD tests with either stationary or tilted kernels are not quantitatively robust to any neighborhoods strictly larger than the singleton set $\{P\}$. We provide numerical evidence in Section˜5.1.